calculate the number of grams of aluminum sulfate produce if 3.85 g of aluminum is added to 6.85 g of sulfuric acid.
1
Expert's answer
2021-10-18T03:09:24-0400

2Al + 3H2SO4 → Al2(SO4)3 + 3H2

Moles of aluminum = 3.85 ÷ 26.99 =0.42 moles


Moles of sulfuric acid = 6.85 ÷98.079 = 0.07moles


Molar mass of aluminium sulfate : 342.15 g/mol


Mass = "342.15 × \frac{0.42}{2}=71.85grams"


Mass"=342.15 × \frac{0.07}{3}=7.9835grams"